### v1

Abstract
Controllable molecular graph generation is essential for material and drug discovery, where generated molecules must satisfy diverse property constraints. While recent advances in graph diffusion models have improved generation quality, their effectiveness in multi-conditional settings remains limited due to reliance on joint conditioning or continuous relaxations that compromise fidelity. To address these limitations, we propose Composable Score-based Graph Diffusion model (CSGD), the first model that extends score matching to discrete graphs via concrete scores, enabling flexible and principled manipulation of conditional guidance. Building on this foundation, we introduce two score-based techniques: Composable Guidance (CoG), which allows fine-grained control over arbitrary subsets of conditions during sampling, and Probability Calibration (PC), which adjusts estimated transition probabilities to mitigate train–test mismatches. Empirical results on four molecular datasets show that CSGD achieves state-of-the-art performance, with a 15.3% average improvement in controllability over prior methods, while maintaining high validity and distributional fidelity. Our findings highlight the practical advantages of score-based modeling for discrete graph generation and its capacity for flexible, multi-property molecular design.
1 Introduction
Molecular graph generation is pivotal for material and drug discovery, enabling efficient exploration of vast chemical spaces [Wang et al.(2025)Wang, Chen, Ma, Yu, Wang, Liu, Ye, Sakurai, and Zeng, Du et al.(2024)Du, Jamasb, Guo, Fu, Harris, Wang, Duan, Liò, Schwaller, and Blundell]. For real-world applications, a key challenge lies in controlling the generation process so that molecules satisfy multiple desired properties [Gebauer et al.(2022)Gebauer, Gastegger, Hessmann, Müller, and Schütt]. For instance, one may wish to generate a set of candidate molecules that simultaneously meet specific ranges of synthetic scores and permeability to target gases.
Recent advances in graph diffusion models [Vignac et al.(2023)Vignac, Krawczuk, Siraudin, Wang, Cevher, and Frossard, Xu et al.(2024)Xu, Qiu, Chen, Chen, Fan, Pan, Zeng, Das, and Tong, Liu et al.(2024)Liu, Xu, Luo, and Jiang] have significantly improved molecular generation quality. However, their ability in multi-conditional generation remains to be improved. Early approaches [Bilodeau et al.(2022)Bilodeau, Jin, Jaakkola, Barzilay, and Jensen, Lee et al.(2023)Lee, Jo, and Hwang] collapse all properties into a single joint condition (Figure 1A), which obscures the influence of individual properties and hinders fine-grained control. GraphDiT [Liu et al.(2024)Liu, Xu, Luo, and Jiang] mitigates this issue by designing separate encoders for categorical and numerical conditions (Figure 1C), yet still treats the combination of encodings as a monolithic condition. This rigid design prevents dynamic inclusion or exclusion of conditions.
Figure 1: Comparison of multi-conditional guidance in discrete graph diffusion: (A) vanilla conditional generation; (B) predictor guidance in DiGress; (C) predictor-free guidance in GraphDiT; (D) our composable score-based guidance. For clarity, the sampling process and notation are simplified.
Therefore, a natural question arises: Is it viable to flexibly compose the guidance of each property considered for molecular graph generation? In analogy to classifier guidance in continuous domain [Dhariwal and Nichol(2021)], DiGress [Vignac et al.(2023)Vignac, Krawczuk, Siraudin, Wang, Cevher, and Frossard] proposes predictor-guidance by combining the gradients from multiple property predictors (Figure 1B). However, this requires relaxing discrete graphs into continuous variables, potentially degrading generation fidelity. Meanwhile, existing discrete graph diffusion models [Vignac et al.(2023)Vignac, Krawczuk, Siraudin, Wang, Cevher, and Frossard, Xu et al.(2024)Xu, Qiu, Chen, Chen, Fan, Pan, Zeng, Das, and Tong, Siraudin et al.(2024)Siraudin, Malliaros, and Morris] typically adopt the mean prediction [Austin et al.(2021)Austin, Johnson, Ho, Tarlow, and Van Den Berg, Campbell et al.(2022)Campbell, Benton, De Bortoli, Rainforth, Deligiannidis, and Doucet], where the denoiser predicts the clean sample’s distribution based on the current noisy and the given condition, rather than the score of the current marginal distribution. This design makes score-based compositional guidance inapplicable [Liu et al.(2022)Liu, Li, Du, Torralba, and Tenenbaum], and directly composing the estimated distributions of each considered condition lacks theoretical interpretation.
To overcome these limitations, we propose Composable Score-based Graph Diffusion model (CSGD), a novel model for flexible multi-conditional molecular generation. Built upon the recently proposed concrete score [Lou et al.(2024)Lou, Meng, and Ermon], we introduce concrete score-based diffusion model for discrete graphs, extending score matching to this domain. This design unlocks advanced score manipulation techniques. Specifically, we introduce Composable Guidance, a novel mechanism that leverages the compositional nature of the score for multi-conditional generation (Figure 1D). Then, we explore Probability Calibration, manipulating conditional concrete scores to better shape the estimated transition probabilities within the generation process. With these techniques, CSGD achieves flexible, property-aware sampling and significantly improves both validity and controllability. In summary, our contributions are as follows:
  • To the best of our knowledge, we introduce the first concrete score-based graph diffusion model, which extends score matching to discrete graph generation and enabling advanced score-manipulation techniques that were previously inapplicable.
  • We propose Composable Guidance and Probability Calibration for multi-conditional molecular generation, which manipulate per-property scores to steer generation toward satisfying desired properties, thereby improving the validity and controllability of the generated molecules.
  • We compare CSGD with various multi-conditional molecular generation methods (Figure 1A-C) on four public datasets. CSGD achieves state-of-the-art performance, with an average 15.3% improvement in controllability, highlighting the effectiveness of score-based modeling and composable guidance.
2 Related Works
Graph Diffusion Models. Diffusion models [Ho et al.(2020)Ho, Jain, and Abbeel, Austin et al.(2021)Austin, Johnson, Ho, Tarlow, and Van Den Berg, Song et al.(2021)Song, Sohl-Dickstein, Kingma, Kumar, Ermon, and Poole] have been widely applied to graph generation. Early approaches such as EDP-GNN [Niu et al.(2020)Niu, Song, Song, Zhao, Grover, and Ermon] and GDSS [Jo et al.(2022)Jo, Lee, and Hwang] treated graphs as continuous variables, typically using a Gaussian reference distribution, similar to that in image generation. Because graphs are inherently discrete, subsequent work [Vignac et al.(2023)Vignac, Krawczuk, Siraudin, Wang, Cevher, and Frossard, Liu et al.(2024)Liu, Xu, Luo, and Jiang, Siraudin et al.(2024)Siraudin, Malliaros, and Morris] modeled graphs in discrete state spaces to better capture their structural properties. Among them, DiGress [Vignac et al.(2023)Vignac, Krawczuk, Siraudin, Wang, Cevher, and Frossard] demonstrated strong performance on various graph generation benchmarks and thus has been widely adopted in practice. However, these models generally operate in a discrete time setting, which limits the flexibility of sampling [Xu et al.(2024)Xu, Qiu, Chen, Chen, Fan, Pan, Zeng, Das, and Tong]. To address this, recent methods such as DisCo [Xu et al.(2024)Xu, Qiu, Chen, Chen, Fan, Pan, Zeng, Das, and Tong] and Cometh [Siraudin et al.(2024)Siraudin, Malliaros, and Morris] first introduce discrete-state, continuous-time diffusion models for graph generation. In this work, our proposed CSGD also operates with continuous time and discrete state, while distinct from previous methods by introducing the concrete score [Lou et al.(2024)Lou, Meng, and Ermon], thereby unlocking score-manipulation techniques for guided sampling that would otherwise be unavailable for the existing mean prediction-based graph diffusion models.
Guidance for diffusion model. Recent studies [Hu et al.(2023)Hu, Zhang, Asano, Burghouts, and Snoek, Ahn et al.(2024)Ahn, Cho, Min, Jang, Kim, Kim, Park, Jin, and Kim, Liu et al.(2022)Liu, Li, Du, Torralba, and Tenenbaum] have shown that sampling guidance techniques, such as classifier guidance [Dhariwal and Nichol(2021)] and classifier-free guidance [Ho and Salimans(2021)], are crucial for steering diffusion models toward samples that better respect satisfy condition(s). As classifier-free guidance has shown its superiority in many tasks, many recent works proposed improvements to it, including applying guidance only during the middle interval of sampling to improve inference speed and quality [Kynkäänniemi et al.(2024)Kynkäänniemi, Aittala, Karras, Laine, Aila, and Lehtinen], guiding a high-quality model with a degraded, lower-quality model trained on the same task and data [Karras et al.(2024)Karras, Aittala, Kynkäänniemi, Lehtinen, Aila, and Laine], a first-principles nonlinear correction [Zheng and Lan(2024)], sliding window guidance [Kaiser et al.(2024)Kaiser, Adaloglou, and Kollmann], and independent condition guidance [Sadat et al.(2025)Sadat, Kansy, Hilliges, and Weber]. These techniques have proven effective in improving the sample quality and condition obedience. Building on our introduction of concrete score, we further propose two score-manipulation techniques to improve the validity and controllability of discrete graph generation.
3 Methodology
In this section, we present CSGD, which tailors SEDD [Lou et al.(2024)Lou, Meng, and Ermon] to the domain of multi-conditional molecular generation (see Section 3.1). Leveraging the composable nature of the scores, CSGD enables more flexible and controllable sampling strategies for generation tasks conditioning on multiple properties (see Section 3.2). In analogy to recently proposed tricks for scores in the continuous domain, we propose Probability Calibration (PC) mechanism for CSGD to improve the validity and controllability of the generated molecular graphs (see Section 3.3).
Notations and Problem Definition. We model graphs in a discrete space, where nodes and edges take categorical attributes, denoted by $\mathcal{X}=\{1,\ldots,a\}$ and $\mathcal{E}=\{1,\ldots,b\}$, respectively. A graph is represented as $G=(\mathbf{X},\mathbf{E})$, where $\mathbf{X}=(x^{i})_{i=1}^{n}\in\mathcal{X}^{n}$ specifies each node’s type (i.e., atom type) and $\mathbf{E}=(e^{ij})_{i,j=1}^{n}\in\mathcal{E}^{n\times n}$ specifies each edge’s type (i.e., bond type or the indicator of no connection). Here, $n$ denotes the number of nodes. Let $\mathcal{C}=\{c_{1},c_{2},\dots,c_{M}\}$ denote a set of $M$ specified conditions, where $c_{i}$ may be numerical or categorical. The multi-conditional molecular generation is to model: $p(G|\mathcal{C})=p(G|c_{1},\dots,c_{M})\propto p(G)p(c_{1},\dots,c_{M}|G)$, where molecular graph $G$ can be evaluated along two dimensions: (1) Distribution Learning $p(G)$: measuring the fidelity to the underlying unconditional distribution; (2) Controllability $p(c_{1},\dots,c_{M}|G)$: measuring consistency between generated molecular graph and the specified conditions. Moreover, there are often practical scenarios in which we need to sample $G$ with only a subset of properties specified. A flexible solution is demanded that can handle an arbitrary subset of properties based on one common model.
3.1 Score-based Graph Diffusion Model
For brevity, we present our diffusion model in an unconditional setting to focus on the use of the concrete score [Meng et al.(2022)Meng, Choi, Song, and Ermon]. This unconditional formulation can be readily adapted to our target multi-conditional generation problem simply by feeding the encoded condition(s) into the denoiser.
Diffusion Process. Following prior graph diffusion models [Xu et al.(2024)Xu, Qiu, Chen, Chen, Fan, Pan, Zeng, Das, and Tong], we diffuse each node and each edge independently following a continuous-time Markov process, which can be described by linear ODEs [Lou et al.(2024)Lou, Meng, and Ermon]:
$\frac{\mathrm{d}p_{t}^{X}}{\mathrm{d}t}=Q_{t}^{X}p_{t}^{X},\quad\frac{\mathrm{d}p_{t}^{E}}{\mathrm{d}t}=Q_{t}^{E}p_{t}^{E},$ (1)
where $\forall t\in\Re_{+},p_{t}^{X}\in\Re^{a},p_{t}^{E}\in\Re^{b}$ denote the marginal discrete probability distributions, with boundary conditions $p_{0}\approx p_{\text{data}}$ and $p_{t\to\infty}\approx p_{\text{base}}$. With a scalar noise schedule $\sigma(t)$ [Ho et al.(2020)Ho, Jain, and Abbeel], the time-dependent transition matrices are defined as $Q_{t}^{X}=\sigma(t)Q^{X}$ and $Q_{t}^{E}=\sigma(t)Q^{E}$, where $Q^{X}\in\mathbb{R}^{a\times a}$ and $Q^{E}\in\mathbb{R}^{b\times b}$ are the transition matrices for nodes and edges, respectively. Due to this independence, the likelihood of a graph $G_{t}$ at timestep $t$ is factorized as $p_{t}(G_{t})=\prod_{i=1}^{n}p_{t}^{X}(x_{t}^{i})\prod_{i,j=1}^{n}p_{t}^{E}(e_{t}^{ij})$. We adopt two common choices for the transition matrix $Q^{X}$ (resp. $Q^{E}$): Absorb [Campbell et al.(2022)Campbell, Benton, De Bortoli, Rainforth, Deligiannidis, and Doucet], which introduces a dedicated absorbing (MASK) state, and Uniform [Austin et al.(2021)Austin, Johnson, Ho, Tarlow, and Van Den Berg], which assigns the same transition probabilities to all other states.
Then, the element-wise forward transition admits a closed form. Conditioned on the state at $t=0$, the probability is given by the cumulative transition matrix:
$p_{t|0}^{X}(\cdot|x^{i}_{0}=x)$ $=x\text{-th column of}\exp{(\bar{\sigma}(t)Q^{X})},$ (2)
$p_{t|0}^{E}(\cdot|e^{ij}_{0}=e)$ $=e\text{-th column of}\exp{(\bar{\sigma}(t)Q^{E})},$
where $\bar{\sigma}(t)=\int_{0}^{t}\sigma(s)\,\mathrm{d}s$ is the cumulative noise coefficient. Thus, the likelihood of the entire graph $G_{t}$ conditioned on the observation $G_{0}$ is given by $p_{t|0}(G_{t}|G_{0})=\prod_{i=1}^{n}p_{t|0}^{X}(x_{t}^{i}|x_{0}^{i})\;\prod_{i,j=1}^{n}p_{t|0}^{E}(e_{t}^{ij}|e_{0}^{ij})$.
Reverse Process. The reverse process aims to reconstruct graphs obeying $p_{\text{data}}$ from noisy ones sampled from $p_{\text{base}}$, which can be completed via simulating the reversal of Eq. (1), that is, $\frac{\mathrm{d}p_{T-t}^{X}}{\mathrm{d}t}=\bar{Q}_{T-t}^{X}p_{T-t}^{X}$ and $\frac{\mathrm{d}p_{T-t}^{E}}{\mathrm{d}t}=\bar{Q}_{T-t}^{E}p_{T-t}^{E}$ [Kelly(1981), Sun et al.(2023)Sun, Yu, Dai, Schuurmans, and Dai]. Since the flux should be consistent, that is, $\bar{Q}_{t}^{X}(x^{\prime},x)p_{t}^{X}(x)=Q_{t}^{X}(x,x^{\prime})p_{t}^{X}(x^{\prime})$, existing works [Meng et al.(2022)Meng, Choi, Song, and Ermon] have proposed to estimate the ratio $\frac{p_{t}^{X}(x^{\prime})}{p_{t}^{X}(x)}$, namely concrete score, and use these scores to perform the reverse process. Edges can be treated analogously via $\frac{p_{t}^{E}(e^{\prime})}{p_{t}^{E}(e)}$.
Due to the combinatorial nature of discrete graphs, a general transition matrix defining probabilities between all possible graph pairs is clearly intractable. Thanks to the independence design of our diffusion process, we restrict each transition to flip only a single position—either a node or an edge—at a time. Under this design, the concrete scores can be naturally parameterized using a neural network $s_{\theta}(\cdot,\cdot)$:
$s_{\theta}^{X}(G_{t},t)_{i,\tilde{x}_{t}^{i}}$ $\approx\frac{p_{t}^{X}(x_{t}^{1}\ldots\tilde{x}_{t}^{i}\ldots x_{t}^{n})}{p_{t}^{X}(x_{t}^{1}\ldots x_{t}^{i}\ldots x_{t}^{n})}\text{ for }1\leq i\leq n,\tilde{x}_{t}^{i}\neq x_{t}^{i},$ (3)
$s_{\theta}^{E}(G_{t},t)_{ij,\tilde{e}_{t}^{ij}}$ $\approx\frac{p_{t}^{E}(e_{t}^{11}\ldots\tilde{e}_{t}^{ij}\ldots e_{t}^{nn})}{p_{t}^{E}(e_{t}^{11}\ldots e_{t}^{ij}\ldots e_{t}^{nn})}\text{ for }1\leq i,j\leq n,\tilde{e}_{t}^{ij}\neq e_{t}^{ij}.$
Here, $x_{t}^{i}$ (resp. $e_{t}^{ij}$) denotes the current node (resp. edge) state and $\tilde{x}_{t}^{i}$ (resp. $\tilde{e}_{t}^{ij}$) are possible alternative states from the corresponding discrete space. Specifically, $s_{\theta}(\cdot,\cdot)$ is often a Graph Transformer that processes both node tokens and edge tokens, where the subscripts $(i,\tilde{x}_{t}^{i})$ mean to take its $\tilde{x}_{t}^{i}$-th output dimension at the $i$-th node token, with edges treated analogously.
Following SEDD, we borrow the derivation of Tweedie denoising [Efron(2011)] yet perform $\tau$-leaping, with each token states $x_{s}^{i}$ and $e_{s}^{ij}$ at timestep $s=t-\Delta t,\Delta t\geq 0$ having their conditional probabilities as follows:
$p_{s|t}^{X}(x^{i}_{s}|x^{i}_{t};\theta)$ $=\left(\operatorname{exp}(-\sigma_{t}^{\Delta t}Q^{X})s_{\theta}^{X}(G_{t},t)_{i}\right)_{x^{i}_{s}}\operatorname{exp}(\sigma_{t}^{\Delta t}Q^{X})(x^{i}_{t},x^{i}_{s}),$ (4)
$p_{s|t}^{E}(e^{ij}_{s}|e^{ij}_{t};\theta)$ $=\left(\operatorname{exp}(-\sigma_{t}^{\Delta t}Q^{E})s_{\theta}^{E}(G_{t},t)_{ij}\right)_{e^{ij}_{s}}\operatorname{exp}(\sigma_{t}^{\Delta t})Q^{E})(e^{ij}_{t},e^{ij}_{s}),$
where $\sigma_{t}^{\Delta t}=\overline{\sigma}(t)-\overline{\sigma}(s)$. Specifically, we perform Euler step for each token independently based on Eq. (4) yet with $\tau$-leaping. Essentially, the reverse step factorizes across tokens:
$p_{s|t}(G_{s}|G_{t};\theta)=\prod_{i=1}^{n}p_{s|t}^{X}(x^{i}_{s}|x^{i}_{t};\theta)\prod_{i,j=1}^{n}p_{s|t}(e^{ij}_{s}|e^{ij}_{t};\theta).$ (5)
Optimization. The denoiser $s_{\theta}$ is trained to predict concrete scores by minimizing the Diffusion Weighted Denoising Score Entropy [Lou et al.(2024)Lou, Meng, and Ermon]:
$\mathcal{L}_{t}^{X}(G_{0},G_{t},\theta)=\sigma(t)\sum_{i=1}^{n}\sum_{\tilde{x}^{i}_{t}=1}^{a}(1-\delta_{x^{i}_{t}}(\tilde{x_{t}}^{i}))\left(s_{\theta}^{X}(G_{t},t)_{i,\tilde{x}^{i}}-\frac{p_{t|0}(\tilde{x_{t}}^{i}|x^{i}_{0})}{p_{t|0}(x^{i}_{t}|x^{i}_{0})}\right)\log s_{\theta}^{X}(G_{t},t)_{i,\tilde{x}^{i}},$ (6)
$\mathcal{L}_{t}^{E}(G_{0},G_{t},\theta)=\sigma(t)\sum_{i,j=1}^{n}\sum_{\tilde{e}_{t}^{ij}=1}^{b}(1-\delta_{e^{ij}_{t}}(\tilde{e_{t}}^{ij}))\left(s_{\theta}^{E}(G_{t},t)_{ij,\tilde{e}^{ij}}-\frac{p_{t|0}(\tilde{e_{t}}^{ij}|e^{ij}_{0})}{p_{t|0}(e^{ij}_{t}|e^{ij}_{0})}\right)\log s_{\theta}^{E}(G_{t},t)_{ij,\tilde{e}^{ij}}.$
The Kronecker delta $\delta_{x_{t}^{i}}(\tilde{x}^{i}_{t})$ and $\delta_{e_{t}^{ij}}(\tilde{e}^{ij}_{t})$ ensures that the current state is excluded from the summation. The final loss combines node and edge terms with weight $\lambda$ as $\mathcal{L}_{t}(G_{0},G_{t},\theta)=\mathcal{L}_{t}^{X}(G_{0},G_{t},\theta)+\lambda\mathcal{L}_{t}^{E}(G_{0},G_{t},\theta)$. The score entropy loss is well-suited for recovering the ground-truth concrete scores [Lou et al.(2024)Lou, Meng, and Ermon]. We optimize it using the Adam [Kingma and Ba(2015)] optimizer, with gradient clipping and a learning rate warmup [He et al.(2016)He, Zhang, Ren, and Sun] to stabilize training.
Analysis. Suppose we discretize the reversal of Eq. 1 by timesteps $0=t_{0}<t_{1}<\cdots<t_{T-1}<t_{T}=1$ and simulate the process to sample $G_{0}$, the corresponding estimated distribution is $p(G;\theta)=\sum_{G_{t_{1}},\ldots,G_{t_{T}}}\prod_{k=1}^{T}p_{t_{k-1}|t_{k}}(G_{t_{k-1}}|G_{t_{k}};\theta)p_{\text{base}}(G_{t_{T}})$. Since isomorphic molecular graphs are treated equally, the permutation-invariant $p(G;\theta)$ is preferred.
Proposition 3.1.
The estimated distribution $p(G;\theta)$ is permutation-invariant.
Proof 3.2.
According to the theoretical results of GeoDiff [Xu et al.(2022)Xu, Yu, Song, Shi, Ermon, and Tang], we just need to show that (1) the base distribution $p_{\text{base}}$ is permutation-invariant and (2) the transitions $p_{t_{k-1}|t_{k}}(G_{t_{k-1}}|G_{t_{k}};\theta)$ are permutation-equivariant.
At first, no matter whether to choose Absorb or Uniform, $p_{\text{base}}$ is factorized over i.i.d. node and edge distributions.
Our denoiser is implemented as a Graph Transformer [Liu et al.(2024)Liu, Xu, Luo, and Jiang], which is permutation-equivariant: any permutation $\pi_{n}$ of the input tokens induces the corresponding permutation $\pi_{n}$ on its output tokens. Hence, the transitions $p_{t_{k-1}|t_{k}}(G_{t_{k-1}}|G_{t_{k}};\theta)$ is permutation-equivariant according to Eq. 4.
As these two conditions are satisfied, we have that $p(\pi_{n}.G;\theta)=p(G;\theta)$.
3.2 Multi-conditional Molecular Generation via Concrete Scores
As discussed at the beginning of this Section, we focus on multi-conditional molecular generation, where the observed dataset $\mathcal{D}=\{(G,\mathcal{C})_{l}\}_{l=1}^{N}$ is an unbiased sample drawn from the underlying joint distribution. Our goal is to estimate the conditional distribution $p(G|\mathcal{C})$ so that we can sample molecular graphs that are valid and respect the specified properties $\mathcal{C}$. We will elaborate on how to achieve this goal with our score-based graph diffusion model and further to enable the flexible specification of an arbitrary subset of the $M$ properties.
Classifier-free guidance (CFG). We first employ the same classifier-free guidance (CFG) as GraphDiT (see Figure 1C). During training and inference, all $M$ properties are encoded and combined into a joint condition embedding, which is fed into the backbone of the denoiser to produce multi-conditional score. Then, both this conditional score and the unconditional one are merged to guide generation. Specifically, given a guidance scale $w$, the CFG score is:
$s_{\theta}^{\mathrm{CFG}}(G_{t},\mathcal{C},t)=s_{\theta}(G_{t},\varnothing,t)+w\,\bigl{(}s_{\theta}(G_{t},\mathcal{C},t)-s_{\theta}(G_{t},\varnothing,t)\bigr{)}.$ (7)
As can be seen, CFG treats all properties as a whole $\mathcal{C}$, producing a single monolithic score $s_{\theta}(G_{t},\mathcal{C},t)$ and limiting fine-grained control over individual properties $c_{m}$.
Figure 2: An overview of CSGD with Composable Guidance (CoG) to enable flexible control ($\text{P}.$: Gas Permeability; Synth.: Synthesizability).
Composable Guidance (CoG). We illustrate one simulation step of CSGD with CoG in Figure 2, where the goal is to flexibly generate molecules conditioned on arbitrary subsets of all the $M$ properties. To this end, we rely on the decomposition of multi-conditional probability:
$p(G|c_{1},\ldots,c_{M})\propto p(G)p(c_{1},\ldots,c_{M}|G)=p(G)\prod_{m=1}^{M}p(c_{m}|G)=p(G)\prod_{m=1}^{M}\frac{p(G|c_{m})p(c_{m})}{p(G)},$ (8)
where we assume conditional independence of the properties given $G$, and we substitute each conditional likelihood $p(c_{m}|G)$ by: $p(c_{m}|G)=\frac{p(G|c_{m})p(c_{m})}{p(G)}$. Following prior work on compositional generation [Liu et al.(2022)Liu, Li, Du, Torralba, and Tenenbaum], we adopt this conditional independence assumption to make score composition tractable, while recognizing that handling correlated conditions remains an important direction for future research.
This decomposition leads to the definition of our CoG score [Liu et al.(2022)Liu, Li, Du, Torralba, and Tenenbaum]:
$s_{\theta}^{\mathrm{CoG}}(G_{t},\mathcal{C},t)$ $=s_{\theta}(G_{t},\varnothing,t)\;+\;\sum_{m=1}^{M}w_{m}\Bigl{(}s_{\theta}(G_{t},c_{m},t)-s_{\theta}(G_{t},\varnothing,t)\Bigr{)}.$ (9)
This CoG strategy modulates the contribution of each single property-specific score $s_{\theta}(G_{t},c_{m},t)$ via a corresponding guidance scale $w_{m}$. Such a formulation provides fine-grained control, enabling selective emphasis on specific subsets of properties during sampling.
To enable CoG, we train a the denoiser $s_{\theta}$ following the procedure described in Algorithm 3.2. During training, each condition $c_{m}$ is picked with a fixed probability, $s_{\theta}$ is enabled to jointly learn the conditional score functions $s_{\theta}(G_{t},c_{m},t)$ for each $c_{m}\in\mathcal{C}$ and the unconditional one $s_{\theta}(G_{t},\varnothing,t)$ in a unified way. At inference time (Algorithm 3.2), these scores are combined via the CoG strategy to guide generation toward arbitrary subsets of target properties.
{algorithm}
[tb]
Training a Score-based Graph Diffusion Model for CoG.
[…]
Sample timestep $t\sim\mathcal{U}(0,1)$ and property $c_{m}\in\mathcal{C}$;
Construct noisy graph $G_{t}$: $x^{i}_{t}\sim p_{t|0}^{X}(\cdot|x^{i}_{0})$, $e^{ij}_{t}\sim p_{t|0}^{E}(\cdot|e^{ij}_{0})$ based on Eq. 2;
Randomly set $c_{m}\leftarrow\varnothing$ with a probability to enable unconditional training;
Predict concrete scores: $(s_{\theta}^{X},s_{\theta}^{E})\leftarrow s_{\theta}(G_{t},c_{m},t)$;
Compute loss $\mathcal{L}_{t}(G_{0},G_{t},\theta)=\mathcal{L}_{t}^{X}(G_{0},G_{t},\theta)+\lambda\mathcal{L}_{t}^{E}(G_{0},G_{t},\theta)$ based on Eq. 6;
Update denoiser $s_{\theta}$ via gradients $\nabla_{\theta}\mathcal{L}_{t}$.
{algorithm}
[tb]
Sampling from Score-based Graph Diffusion Model with CoG.
[…]
Sample $G_{T}\sim p_{\text{base}}$, $k\leftarrow T$;
[…]
$k>0$ Compute unconditional score $s_{\theta}(G_{t},\varnothing,t_{k})$;
[…]
each $1\leq l\leq L$Compute conditional score: $s_{\theta}(G_{t},c_{m_{l}},t_{k})$;
Compute CoG scores $s_{\theta}^{\text{CoG}}(G_{t},\mathcal{C}^{\prime},t_{k})$ based on Eq. 9;
Construct transition densities $p_{t_{k-1}|t_{k}}^{X}(x_{t_{k-1}}^{i}|x_{t_{k}}^{i};\theta)$ and $p_{t_{k-1}|t_{k}}^{E}(e_{t_{k-1}}^{ij}|e_{t_{k}}^{ij};\theta)$ by substituting $s_{\theta}$ by $s_{\theta}^{\text{CoG}}$ based on Eq. 4;
Apply PC to transition densities $p_{t_{k-1}|t_{k}}^{X}(x_{t_{k-1}}^{i}|x_{t_{k}}^{i};\theta)$ and $p_{t_{k-1}|t_{k}}^{E}(e_{t_{k-1}}^{ij}|e_{t_{k}}^{ij};\theta)$ as described in Section $3\tcp*$Optional
[…]
Construct $G_{t_{k-1}}\leftarrow((x_{t_{k-1}}^{i})_{i=1}^{n},(e_{t_{k-1}}^{ij})_{i,j=1}^{n})$;
[…]
3.3 Probability Calibration with Concrete Scores
After obtaining the guided score, we compute the transition probabilities $p^{\theta}_{s|t}(G_{s}|G_{t};\theta)$ according to Eq. (5). Although guidance improves condition alignment, it may also compromise sample fidelity. In particular, high guidance scales can cause a train–test mismatch [Saharia et al.(2022)Saharia, Chan, Saxena, Li, Whang, Denton, Ghasemipour, Gontijo Lopes, Karagol Ayan, Salimans, et al.], where the guided score, i.e., $s_{\theta}^{\mathrm{CoG}}$, during sampling falls outside the ranges observed in training. In our work, we observe a similar phenomenon which degrades both the validity and controllability of generated molecules. To address this, we apply dynamic thresholding and temperature scaling, which jointly calibrate the probabilities.
Dynamic Thresholding. Let $p_{\text{raw}}$ denote the raw guided probabilities computed with $s_{\theta}^{\text{CoG}}$ based on Eq. 4. We first clamp negative entries to zero and compute the lower and upper percentiles (e.g., $\alpha=1\%$, $\beta=99\%$) across the state dimension: $\ell=\mathrm{Quantile}(p_{\mathrm{raw}},\alpha),h=\max\bigl{(}\mathrm{Quantile}(p_{\mathrm{raw}},\beta),\,\ell+\varepsilon\bigr{)},$ where $\varepsilon$ is a small constant (e.g., $10^{-6}$) to avoid division by zero. The probabilities are then normalized by min–max scaling:
$p_{\mathrm{clipped}}=\frac{\bigl{[}p_{\mathrm{raw}}-\ell\bigr{]}_{+}}{\,h-\ell\,},\qquad[p]_{+}\equiv\max(p,0).$ (10)
This procedure guarantees valid probabilities in $[0,1]$, stabilizing sampling by preventing extreme values from dominating. Our dynamic thresholding serves as a discrete analogue to Imagen’s method [Saharia et al.(2022)Saharia, Chan, Saxena, Li, Whang, Denton, Ghasemipour, Gontijo Lopes, Karagol Ayan, Salimans, et al.]: rather than scaling continuous pixel values, we calibrate discrete transition probabilities by percentile-based clipping, ensuring well regularized probability ranges during sampling.
Temperature Scaling. Temperature scaling is a post-processing method used to improve probabilistic predictions in classification tasks [Guo et al.(2017)Guo, Pleiss, Sun, and Weinberger, Kull et al.(2019)Kull, Perello Nieto, Kängsepp, Silva Filho, Song, and Flach, Xuan et al.(2025)Xuan, Yang, and Li]. Its impact on graph diffusion sampling remains unexplored. Here, we investigate whether temperature scaling can improve the validity and controllability of generated graphs by modulating the sharpness of the predicted distributions at each step of sampling. To this end, we further adjust the distribution via temperature $\tau$:
$p_{\mathrm{scaled}}(i)=\frac{\bigl{(}p_{\mathrm{clipped}}(i)\bigr{)}^{1/\tau}}{\sum_{j}\bigl{(}p_{\mathrm{clipped}}(j)\bigr{)}^{1/\tau}}\text{ for each outcome $i$}.$ (11)
Lower $\tau$ (i.e., $0<\tau<1$) sharpens the distribution and favors high-probability states, while higher $\tau$ (i.e., $\tau>1$) flattens it and promotes diversity.
Probability Calibration combines percentile thresholding and temperature scaling to calibrate guided probability, enhancing both validity and controllability in molecular graph generation. In practice, it involves three key parameters: $\alpha$, $\beta$, and $\tau$. Based on our experience, $\tau$ is the parameter most worth tuning, while $\alpha$ and $\beta$ can be assigned reasonable values to ensure stability (e.g., $\alpha=1\%$, $\beta=99\%$).
4 Experiments
To evaluate the effectiveness of extending score matching to discrete graph diffusion models, as well as the score-based capacities it enables, namely, composable guidance (CoG) and probability calibration (PC), we conducted comprehensive quantitative experiments to answer the following research questions: RQ1: Does CSGD outperform related discrete graph diffusion models? RQ2: Can concrete scores improve the quality and controllability of samples relative to the mean prediction approach? RQ3: How beneficial are CoG and PC for generating valid and controllable samples in multi-conditional generation? RQ4: Can CoG support flexible sampling under varying numbers and types of molecular property conditions?
4.1 Experimental Setup
Datasets. Following the experimental setup in GraphDiT [Liu et al.(2024)Liu, Xu, Luo, and Jiang], we evaluate multi-conditional molecular generation using two types of datasets: (1) For materials design, we use a Polymers dataset [Yampolskii(2012)], which includes four numerical conditions related to synthesizability—Synthetic Accessibility Score (SAS) [Ertl and Schuffenhauer(2009)] and Synthetic Complexity Score (SCS) [Coley et al.(2018)Coley, Rogers, Green, and Jensen]—as well as gas permeability properties ($\text{O}_{2}$P., $\text{N}_{2}$P. and $\text{CO}_{2}$P.). (2) For drug design, we evaluate on three class-balanced dataset–BACE, BBBP, and HIV–from MoleculeNet [Wu et al.(2018)Wu, Ramsundar, Feinberg, Gomes, Geniesse, Pappu, Leswing, and Pande]. Each dataset provides a categorical condition specific to its task (e.g., blood–brain barrier permeability for BBBP, $\beta$-secretase 1 inhibition for BACE, and HIV replication inhibition for HIV), together with the same numerical synthesizability conditions (SAS and SCS) used in the Polymers dataset.
Metrics. To systematically evaluate model performance, we use over three classes of metrics covering generation quality, distribution Learning, and controllability. First, we assess Validity, defined as the proportion of valid molecules among generated samples. Second, we evaluate the model’s ability to learn the target distribution using three metrics: Diversity (measured by Tanimoto similarity [Bajusz et al.(2015)Bajusz, Rácz, and Héberger] among generated molecules), Similarity (fragment-based similarity to the reference set), and FCD [Preuer et al.(2018)Preuer, Renz, Unterthiner, Hochreiter, and Klambauer] (Fréchet ChemNet Distance) between generated and reference distributions. Third, we measure controllability: for numerical conditions, we report the MAE between the specified conditions and predicted properties of the generated molecules; for categorical conditions, we report Accuracy.
Baselines. We compare our method with recent diffusion-based models for multi-conditional molecular generation, including GDSS [Jo et al.(2022)Jo, Lee, and Hwang], DiGress [Vignac et al.(2023)Vignac, Krawczuk, Siraudin, Wang, Cevher, and Frossard], and their conditional variants MOOD [Lee et al.(2023)Lee, Jo, and Hwang] and DiGress v2, as well as GraphDiT [Liu et al.(2024)Liu, Xu, Luo, and Jiang], which represents the current state-of-the-art in this task. For MOOD and DiGress v2, trained multi-task property predictors with the same architecture are used to provide guidance. We note that GraphDiT adopts a graph-dependent forward diffusion process, which demonstrates superior performance within its full model pipeline [Liu et al.(2024)Liu, Xu, Luo, and Jiang]. This creates a noise-model mismatch relative to our independent tokenwise scheme. If anything, this discrepancy may favor GraphDiT, making our evaluation of CSGD conservative and thus fair under these conditions.
Implementation Details. Our implementation builds on the official open-source code of GraphDiT [Liu et al.(2024)Liu, Xu, Luo, and Jiang], leveraging its Graph Transformer and condition encoder to handle both numerical and categorical properties. We split each dataset into training, validation, and test sets with a 6:2:2 ratio. For evaluation, we generate 10,000 molecules conditioned on the properties of the reference set. Generation validity and distributional metrics are computed using the MOSES [Polykovskiy et al.(2020)Polykovskiy, Zhebrak, Sanchez-Lengeling, Golovanov, Tatanov, Belyaev, Kurbanov, Artamonov, Aladinskiy, Veselov, et al.]. For controllability evaluation, we follow GraphDiT and adopt a random forest model trained on all task-related molecules as an Oracle [Gao et al.(2022)Gao, Fu, Sun, and Coley]. The properties of samples are then compared with the specified conditions to calculate MAE/Accuracy for numerical/categorical property.
Table 1: Results on the Polymers dataset including four numerical conditions. The best-performing methods for each metric are highlighted in bold, and the second-best are underlined.
[…]
Table 2: Comparisons with both numerical and categorical conditions evaluated by MAE and Accuracy, respectively. The best-performing methods for each metric are highlighted in bold, and the second-best are underlined.
[…]
4.2 (RQ1) Overall Performance Comparison
We first compare the overall performance of CSGD with both CoG and PC against existing baselines across all the adopted datasets.
We present the results on the Polymers dataset in Table 1, where CSGD consistently achieves notable improvements over the most competing baseline GraphDiT in both validity and controllability. With a transition matrix leading to the Uniform $p_{\text{base}}$, CSGD attains an average relative improvement in MAE of +15.3% compared to the strongest baseline. While DiGress and its v2 variant generate diverse molecules, they struggle to satisfy multiple conditions simultaneously, leading to weaker controllability and lower FCD scores. In contrast, CSGD not only maintains strong distribution learning but also achieves the best FCD, confirming the effectiveness of CSGD.
Results on datasets with both numerical and categorical conditions (see Table 2) further confirm these findings. CSGD achieves the best validity and controllability, while maintaining competitive distribution learning, particularly in terms of FCD and similarity. These results demonstrate that CSGD can flexibly guide the generation of molecular graphs in diverse conditional settings, producing valid molecules with improved compliance to the specified properties.
4.3 Ablation Studies
To assess the contribution of each component, we conduct a detailed ablation study focusing on two aspects: score-based modeling in graph diffusion and two score manipulation techniques.
4.3.1 (RQ2) Performance of Score-based Modeling
Score-based modeling is evaluated against the current mean-prediction approach, GraphDiT, using the same guidance strategy. Due to space constraints, we report results on two datasets in Table 3, with the complete results across all four datasets provided in Appendix (see Table 6). For clarity, results on the Polymers dataset are reported as the average MAE of gas permeability across all gases, rather than per-gas MAE.
The most notable improvement is in molecular validity, with CSGD achieving better performance under both Absorb and Uniform $p_{\text{base}}$. Improvements in controllability also remain competitive. Taken together, these findings indicate that score-based modeling substantially improves the validity and controllability of discrete graph diffusion in multi-conditional generation.
Table 3: Ablation (RQ2): Comparing CSGD (w/o PC, and with CFG rather than CoG) to GraphDiT with CFG, where numerical and categorical conditions are evaluated using MAE and Accuracy, respectively.
[…]
Table 4: Ablation (RQ3): Comparison of CoG and CFG in CSGD with and without PC. Numerical and categorical conditions are evaluated using MAE and Accuracy, respectively.
[…]
4.3.2 (RQ3) Effectiveness of Composable Guidance and Probability Calibration
As shown in Table 4, we evaluate the effectiveness of CoG on two datasets, with results on all four benchmarks provided in the Appendix (see Table 7). Under the same network architecture, hyperparameters, and guidance scale, CoG achieves performance comparable to CFG with a Uniform $p_{\text{base}}$. For instance, CSGD with CoG attains 98.84% validity while maintaining strong controllability. Across other datasets, CoG consistently matches CFG in performance while enabling flexible sampling of CSGD, underscoring its practicality for multi-conditional molecular generation.
We further assess the effect of PC under both CFG and CoG. Across all datasets and guidance strategies, incorporating PC consistently enhances validity and controllability. Despite its simplicity, PC proves to be a practical and effective tool, serving as a general, plug-and-play component to improve diverse score-manipulation techniques in molecular graph generation.
Table 5: (RQ4): Comparison of CSGD performance when sampling with different subsets of properties specified. Total guidance weight of each subset is kept the same. Synth.: Synthesizability; $\text{O}_{2}$, $\text{N}_{2}$, $\text{CO}_{2}$: Gas Permeability.
Model
Property
$\text{Subsets}^{*}$
[…]
GraphDiT {Synth., $\text{O}_{2}$, $\text{N}_{2}$, $\text{CO}_{2}$} 82.45 1.2973 0.7440 0.8857 0.7550
CSGD {Synth., $\text{O}_{2}$, $\text{N}_{2}$, $\text{CO}_{2}$} 99.38 1.1905 0.6054 0.6993 0.6232
{Synth.} 99.78 0.9179 - - -
$\text{O}_{2}$, $\text{N}_{2}$, $\text{CO}_{2}$} 99.38 - 0.6001 0.6927 0.6133
{Synth., $\text{O}_{2}$} 99.50 1.1894 0.6099 - -
4.4 (RQ4) Flexibility
To evaluate the flexibility of CSGD, we conduct experiments on the Polymers dataset, which contains two types of four properties and is well-suited for this analysis. Using the same trained models, we assess CSGD under different property subsets, including the full set, single-type subsets (e.g., Synthesizability alone or only Gas Permeability properties), and mixed-type subsets (e.g., Synthesizability combined with a single Gas Permeability property).
As shown in Table 5, CoG enables flexible sampling by conditioning generation on arbitrary subsets of properties. When conditioning solely on Synthesizability or the Gas Permeability properties ($\text{O}_{2}$, $\text{N}_{2}$, $\text{CO}_{2}$), the model achieves the best controllability for the targeted property while maintaining high validity, suggesting that focusing on a single property allows more precise control. For mixed subsets (e.g., Synthesizability combined with $\text{O}_{2}$), the model effectively balances multiple objectives, achieving strong validity and good controllability on the selected properties. Overall, these results demonstrate that CoG supports flexible and practical multi-conditional generation, enabling molecular design to be tailored to diverse application needs.
4.5 Conclusion and Future Directions
In this work, inspired by the flexibility of score-based modeling in continuous-state diffusion, we introduce concrete scores for discrete graph diffusion. By bridging the gap between discrete molecular structures and score-based modeling, concrete scores open the door to applying a wide range of score-based techniques to discrete graph domains. To demonstrate their flexibility and effectiveness, we propose two novel strategies and validate them on the challenging task of multi-conditional molecular generation. We demonstrate that CSGD achieves strong performance and exhibits high flexibility on several multi-conditional molecular generation benchmarks.
Recent progress in discrete flow matching—notably Discrete Flow Matching (DFM) for general discrete data [Gat et al.(2024)Gat, Remez, Shaul, Kreuk, Chen, Synnaeve, Adi, and Lipman] and its graph-specialized variant DeFoG [Qin et al.(2025)Qin, Madeira, Thanou, and Frossard]—offers an alternative generative paradigm that disentangles training and sampling for increased efficiency. In contrast, our method operates within a score-based diffusion framework, enabling principled score composability and flexible multi-conditional control; exploring hybrid approaches that combine composable score guidance with the sampling efficiency of discrete flow matching is an exciting direction for future work. Moreover, we hope to extend our framework to a wider range of graph generation domains such as citation and social networks, providing a unified foundation for controllable graph generative modeling.

### v2

@@
 1 Introduction
-Molecular graph generation is pivotal for material and drug discovery, enabling efficient exploration of vast chemical spaces [Wang et al.(2025)Wang, Chen, Ma, Yu, Wang, Liu, Ye, Sakurai, and Zeng, Du et al.(2024)Du, Jamasb, Guo, Fu, Harris, Wang, Duan, Liò, Schwaller, and Blundell].
-For real-world applications, a key challenge lies in controlling the generation process so that molecules satisfy multiple desired properties [Gebauer et al.(2022)Gebauer, Gastegger, Hessmann, Müller, and Schütt].
+Molecular graph generation is pivotal for material and drug discovery, enabling efficient exploration of vast chemical spaces [Wang et al.(2025)Wang, Chen, Ma, Yu, Wang, Liu, Ye, Sakurai, and Zeng, Du et al.(2024)Du, Jamasb, Guo, Fu, Harris, Wang, Duan, Liò, Schwaller, and Blundell, Huang et al.(2024)Huang, Sun, Du, and Lv].
+For real-world applications, a key challenge lies in controlling the generation process so that molecules satisfy multiple desired properties [Gebauer et al.(2022)Gebauer, Gastegger, Hessmann, Müller, and Schütt, Zhang et al.(2025)Zhang, Huang, Lou, Ding, Cao, and Wang].
 For instance, one may wish to generate a set of candidate molecules that simultaneously meet specific ranges of synthetic scores and permeability to target gases.
@@
 We will elaborate on how to achieve this goal with our score-based graph diffusion model and further to enable the flexible specification of an arbitrary subset of the $M$ properties.
+Condition Encoders.
+To handle multi-conditional generation tasks that involve both categorical and numerical properties, we adopt separate encoders for each property, following GraphDiT.
+Specifically, each categorical property is processed by a Categorical Encoder, and each numerical property by a Cluster Encoder.
+See Appendix 7.0.3 for details of these encoders.
+The resulting embeddings can then be combined in different ways, such as aggregating all property embeddings to form a joint representation with the mean pooling (see Figure 1C) or using a single property embedding as the input to the denoiser (see Figure 1D).
 Classifier-free guidance (CFG).
@@
 As can be seen, CFG treats all properties as a whole $\mathcal{C}$, producing a single monolithic score $s_{\theta}(G_{t},\mathcal{C},t)$ and limiting fine-grained control over individual properties $c_{m}$.
-Figure 2: An overview of CSGD with Composable Guidance (CoG) to enable flexible control ($\text{P}.$: Gas Permeability; Synth.: Synthesizability).
+Figure 2: An overview of CSGD with Composable Guidance (CoG) to enable flexible control ($\text{P}.$: Gas Permeability; Synth.: Synthesizability). 1 Dynamic Thresholding, 2 Temperature Scaling.
 Composable Guidance (CoG).
@@
 At inference time (Algorithm 3.2), these scores are combined via the CoG strategy to guide generation toward arbitrary subsets of target properties.
+Efficiency Concern With CoG, a separate score must be computed for each property, that is to say, the inference time increase linearly with the number of specified properties, which would be costly when this number is large.
+To eliminate this issue, we propose a training strategy that enables both CoG and a more efficient variant, termed fast-CoG.
+During training, instead of uniformly sampling one property $c_{m}$ (Line3 in Algorithm 3.2), we randomly sample an non-empty subset $\mathcal{C}_{\text{sub}}\subseteq\mathcal{C}$ of the $M$ specified properties and aggregate their embeddings via mean pooling to form a single combined condition, which will be fed into the denoiser.
+The model is trained to predict the corresponding multi-conditional score $s_{\theta}(G_{t},\mathcal{C}_{\text{sub}},t)$ based on this aggregated embedding.
+At inference time, $s_{\theta}$, learned in this way, enables us to compute the desired multi-conditional score for any subset of specified properties with one time of model inference, which can then be directly used to make state transition.
+Formally, the fast-CoG score is defined as:
+$s_{\theta}^{\mathrm{fast-CoG}}(G_{t},\mathcal{C}_{\text{sub}},t)=s_{\theta}(G_{t},\varnothing,t)+w\,\bigl(s_{\theta}(G_{t},\mathcal{C}_{\text{sub}},t)-s_{\theta}(G_{t},\varnothing,t)\bigr).$ (10)
+Unlike CFG, fast-CoG allows flexible selection of property subsets during sampling.
+This training strategy enables the model to support both CoG and fast-CoG at inference, providing an option for scenarios that prioritize efficiency.
 {algorithm}
@@
 […]
-Sample timestep $t\sim\mathcal{U}(0,1)$ and property $c_{m}\in\mathcal{C}$;
+Sample timestep $t\sim\mathcal{U}(0,1)$ and property $m\sim\mathcal{U}(\{1,\ldots,M\})$;
 Construct noisy graph $G_{t}$: $x^{i}_{t}\sim p_{t|0}^{X}(\cdot|x^{i}_{0})$, $e^{ij}_{t}\sim p_{t|0}^{E}(\cdot|e^{ij}_{0})$ based on Eq. 2;
@@
 RQ4: Can CoG support flexible sampling under varying numbers and types of molecular property conditions?
+Then, we assess the sampling efficiency of CSGD with various guidance strategies to show its versatility and flexibility.
 4.1 Experimental Setup
@@
 Overall, these results demonstrate that CoG supports flexible and practical multi-conditional generation, enabling molecular design to be tailored to diverse application needs.
-4.5 Conclusion and Future Directions
+4.5 Efficiency
+In the above experiments, we evaluate the performance and flexibility of CSGD using CoG.
+However, a natural concern arises: the inference time of CoG grows linearly with the number of specified properties, which may limit its practicality in efficiency-sensitive scenarios.
+To address this, in Sec.
+3.2, we introduce a training strategy that enables CSGD to support both CoG and a more efficient variant, fast-CoG, simultaneously.
+As shown in Table 6, we measure the average inference time per molecule on the test set using CSGD trained with that training strategy, across different property subsets and guidance strategies.
+The results confirm that the inference time of CoG is longer.
+In contrast, fast-CoG achieves comparable performance while maintaining more efficient sampling.
+This demonstrates the strong flexibility of CSGD: a single trained model can switch between CoG and fast-CoG with different property subsets at inference, allowing users to balance controllability and efficiency according to the needs of different scenarios.
+Table 6: Evaluation of inference efficiency of CSGD using CoG and fast-CoG across different subsets of properties specified.
+All results are derived from the same trained model, using different sampling settings at inference.
+denotes the inference time per molecule, measured in seconds.
+[…]
+5 Conclusion and Future Directions
 In this work, inspired by the flexibility of score-based modeling in continuous-state diffusion, we introduce concrete scores for discrete graph diffusion.